OpenStudy (anonymous):

What is the resistance of a 7.4-m length of copper wire 1.2mm in diameter? The resistivity of copper is 1.68×10−8Ω⋅m.

hartnn (hartnn):

know the formula?

OpenStudy (anonymous):

is it R = V/A

OpenStudy (anonymous):

I dont think its that one...

hartnn (hartnn):

\(\Large R = \rho \dfrac{L}{A}\)

hartnn (hartnn):

\(\rho \) is resistivity L is Length A is the area

hartnn (hartnn):

since diameter is given, first you need to find Area can you ?

OpenStudy (anonymous):

do I use the formula for a cylinder?

OpenStudy (anonymous):

A = 2 pi (1.2mm) + 2 pi (1.2mm) (7.4m)

OpenStudy (anonymous):

i got 46.5

hartnn (hartnn):

its the cross-sectional area only A = \(\pi r^2 = \pi d^2/4\)

OpenStudy (anonymous):

ohh so it is 1.13x 10^-6 =A

hartnn (hartnn):

yes good :)

hartnn (hartnn):

so now you know all, \(\rho,L,A\)

OpenStudy (anonymous):

for an answer i got 0.1100176991

OpenStudy (anonymous):

R = (1.68x10^-8) (7.4 / 1.13x10^-6)

hartnn (hartnn):

seems to be correct :)

OpenStudy (anonymous):

is the unit ohm

hartnn (hartnn):

yes, unit of resistance is ohm

OpenStudy (anonymous):

thank you very much

hartnn (hartnn):

welcome ^_^